OpenStudy (anonymous):

How do I solve -4√(7)√(42), with STEPS included?

OpenStudy (anonymous):

\[\LARGE -4\cdot \sqrt7 \cdot \sqrt{42}=-4\sqrt{7\cdot 42}=...\]

OpenStudy (cwrw238):

= -4*sqrt7*(sqrt7*sqrt6) = -4*7*sqrt6 = -28sqrt6